OpenStudy (anonymous):

Alejandro made an error in the steps below when determining the equation of the line that is perpendicular to the line 4x – 3y = –8 and passes through the point (3, –2). Alejandro made his first error in which step? A Step 1 B Step 2 C Step 3 D Step 4

OpenStudy (mathmale):

Your thoughts? If the slope of a given line is m, what is the slope of a line perpendicular to the given line?

OpenStudy (kayders1997):

If you try the first step do you get what they get?

OpenStudy (kayders1997):

Do you know what per circular slope means?

OpenStudy (kayders1997):

Perpendicular

OpenStudy (anonymous):

now i do

OpenStudy (anonymous):

but i dont know what it means by this problem?

OpenStudy (kayders1997):

should we go through each step until we get to the right answer? Than you'll be able to understand

OpenStudy (anonymous):

yeah

OpenStudy (kayders1997):

Okay so you have the equation 4x-3y=8 in these problems you need to solve for y= so what would you do first?

OpenStudy (anonymous):

plug the 3 in for y

OpenStudy (kayders1997):

Nope move the x part on the other side

OpenStudy (anonymous):

let me get a paper

OpenStudy (anonymous):

ok

OpenStudy (kayders1997):

okay let me know when your ready

OpenStudy (anonymous):

im ready

OpenStudy (kayders1997):

4x-3y=8 like I said you want to get the y's alone on one side so first your going to have to get the x's on the other side do you know how to do that?

OpenStudy (anonymous):

subtract 4 to 8

OpenStudy (kayders1997):

4x don't forget about the x but yes that is the right away so you will have?

OpenStudy (kayders1997):

*idea

OpenStudy (kayders1997):

Do you understand?

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

u mean on the 4x

OpenStudy (anonymous):

because i got that figured out

OpenStudy (kayders1997):

Yes so we're not just subtracting 4 it's 4x is all, so what do you have?

OpenStudy (anonymous):

the 3y left

OpenStudy (kayders1997):

Yes so how would you get rid of the 3 but not the y?

OpenStudy (anonymous):

-3

OpenStudy (kayders1997):

Because we want y=stuff

OpenStudy (anonymous):

-3y

OpenStudy (kayders1997):

Divide? Subtract?

OpenStudy (anonymous):

divide?

OpenStudy (kayders1997):

Yes but not -3 because your have a positive 3 3/3=1 so than you would have 1y or just y left :)

OpenStudy (anonymous):

oh ok

OpenStudy (kayders1997):

So what did you get for the full equation

OpenStudy (kayders1997):

Y=?

OpenStudy (anonymous):

+3y

OpenStudy (anonymous):

im getting confused on the y kind of.

OpenStudy (kayders1997):

Nope we got ride of the 3y because we divided it so so had 4x-3y=8 so we would have subtracted 4x so we have -3y=-4x+8 so yes I'm sorry divide by -3 I thought it was positive 3y so divide by 3 you get y=?

OpenStudy (kayders1997):

Divide by -3

OpenStudy (anonymous):

-1

OpenStudy (kayders1997):

Divide everything by -3 whatever you do to one side you do to the other

OpenStudy (kayders1997):

So -3y/-3 -3/-3y=1y which is y now do that to the other parts

OpenStudy (kayders1997):

*-3y/-3

OpenStudy (mathmale):

If I understand this problem statement correctly, you do NOT have to find the equation of a line perpendicular to the given line. Instead, you are to read through the list of steps taken by this guy Alejandro and identify the step in which he made his FIRST error. That's all. I asked you before: If the slope of the given line is 4/3, then what is the slope of a line perpendicular to the given line? Could you answer this now, please?

OpenStudy (kayders1997):

He wanted to go through the steps

OpenStudy (mathmale):

Please let @Helpplzze speak for himself.

OpenStudy (kayders1997):

Ummmm that's what he said

OpenStudy (mathmale):

Please let @Helpplzze speak for himself.

OpenStudy (kayders1997):

do you understand @Helpplzze

OpenStudy (mathmale):

The key issue in this whole problem is WHAT WOULD BE THE SLOPE OF A LINE PERPENDICULAR TO THE GIVEN LINE?

OpenStudy (kayders1997):

Yes math ale we understand that but he wanted to go through the steps until we found the wrong answer, there's no reason to fight over this it will just give this person more practice anyway

OpenStudy (mathmale):

My concern is that @Helpplzze is not involved in this discussion at all, and you (kayders1997) continue to speak for him. Kindly wind up your discussion. One last time: the key issue here is

OpenStudy (mathmale):

what is the slope of a line perpendicular to the given line? Until @Helpplzze has answered that correctly, all the rest of your discussion is pointless.

OpenStudy (kayders1997):

ummmm really?

OpenStudy (mathmale):

Sorry to bring this up, but @helpplzze has spoken for himself: "Helpplzze Best Response Medals 0 im getting confused on the y kind of." Yes, pointless.

OpenStudy (kayders1997):

It does say step one could be wrong

OpenStudy (mathmale):

And I am suggesting that the way to identify that step is to determine the correct slope of a line perpendicular to the given line.

OpenStudy (mathmale):

Note that our student is now offline.

OpenStudy (kayders1997):

Yeah